3. Determine whether the given value is a root of the function. f(x) = x + 2x2 - 4x - 3; x=-2
Answer #1

# Given that f(x) = x3 + 2x2_4x=3 then we find 25-2 is loot of f(x). we know that 229 root a2+bx+c . fca) al then # tx) = qay